# Steering Away from Harm: An Adaptive Approach to Defending Vision Language Model Against Jailbreaks

### v2

@@ Abstract
Vision Language Models (VLMs) can produce unintended and harmful content when exposed to adversarial attacks, particularly because their vision capabilities create new vulnerabilities. Existing defenses, such as input preprocessing, adversarial training, and response evaluation-based methods, are often impractical for real-world deployment due to their high costs. To address this challenge, we propose ASTRA, an efficient and effective defense by adaptively steering models away from adversarial feature directions to resist VLM attacks. Our key procedures involve finding transferable steering vectors representing the direction of harmful response and applying adaptive activation steering to remove these directions at inference time. To create effective steering vectors, we randomly ablate the visual tokens from the adversarial images and identify those most strongly associated with jailbreaks. These tokens are then used to construct steering vectors. During inference, we perform the adaptive steering method that involves the projection between the steering vectors and calibrated activation, resulting in little performance drops on benign inputs while strongly avoiding harmful outputs under adversarial inputs. Extensive experiments across multiple models and baselines demonstrate our state-of-the-art performance and high efficiency in mitigating jailbreak risks. Additionally, ASTRA exhibits good transferability, defending against both unseen attacks at design time (i.e., structured-based attacks) and adversarial images from diverse distributions. Our code is available at https://github.com/ASTRAL-Group/ASTRA.
@@ 1 Introduction
Jailbreak attacks in VLMs aim to induce models to generate harmful responses by using jailbreaking image-text pairs [20, 19, 46, 43, 37, 22]. These jailbreak attacks can be categorized into two types: (i) perturbation-based attacks, which create adversarial images that prompt generation of the harmful response from VLMs [37, 41, 2, 34], (ii) structured-based attacks, which embeds the malicious queries into images via typography to bypass the safety alignment of VLMs [14, 26]. Countermeasures for both attacks have been explored extensively: the input preprocessing-based method [33] or adversarial training [21] have proven effective for perturbation-based attacks. However, these defenses suffer as they require intensive computations to purify the image or fine-tune the model. Response evaluation-based  [15, 48, 53] defenses have been proposed for structured-based attacks, but they all require running model inference multiple times to potentially identify harmful outputs, which dramatically increases the cost of real-world deployment.
@@
    ASTRA achieves a substantial improvement in defending against perturbation-based attacks. Compared to state-of-the-art methods JailGuard [53], with a Toxicity Score of 12.12% and an Attack Success Rate of 17.84% lower, and 9x faster in MiniGPT-4. ASTRA is also transferable to some unseen attacks (e.g., structure-based attacks) and adversarial images with highly different distributions, and still be effective against adaptive attacks.
@@ 2 Related Work
Jailbreak Attacks on VLM. Jailbreak attacks aim to alter the prompt to trick the model into answering forbidden questions. Apart from the LLM-based textual jailbreak strategies [27, 58, 16, 51], additional visual inputs expose a new attack surface to VLM attacks. There are two main types of attacks: perturbation-based attacks and structured-based attacks [48]. Perturbation-based attacks create adversarial images to bypass the safeguard of VLMs [6, 37, 2, 41, 50, 54]. Structued-based attacks convert the harmful content into images through typography or text-to-image tool (e.g., Stable Diffusion [40]) to induce harmful responses from the model [25, 14, 26, 29, 23] We study our defense on both types of attacks.
@@ 3 Methodology
         Compute: $f(g_{n})=\mathcal{P}_{\text{VLM}}(\mathbf{r}|\text{Ablate}(\mathbf{x}_{v},g_{n}),\mathbf{x}_{t})$
@@ 3.1 Constructing Steering Vectors
Harmful Feature Extraction. With attribution scores for each token, we extract the representation of those tokens strongly correlated with jailbreak. Additionally, we hope our steering vectors are not textual query dependent and enjoy good transferability to a wider range of jailbreaks. Thus, we utilize visual tokens with Top-$k$ attribution scores from the surrogate model $\hat{f}$ paired with the empty user query to construct the steering vectors.
@@
$v^{l}=\frac{1}{|\mathcal{D}|}\sum_{\mathbf{x}_{v},\text{Mask}(\mathbf{x}_{v})\in\mathcal{D}}\mathbf{a}^{l}(\mathbf{x}_{v},\mathbf{x}_{\text{template}})-\mathbf{a}^{l}(\text{Mask}(\mathbf{x}_{v}),\mathbf{x}_{\text{template}})$
@@ 4 Experiments
[…]
SD SD_TYPO TYPO $\epsilon=16/255$ $\epsilon=32/255$ $\epsilon=64/255$ unconstrained
@@
w/o defense 13.75 43.25 43.75 50.91 58.18 56.36 65.45
ASTRA (Ours) 3.75 8.75 11.25 8.18 7.27 6.36 9.09
@@
w/o defense 20.00 61.25 38.75 63.46 63.63 65.45 64.55
ASTRA (Ours) 11.25 40.00 33.75 3.64 3.64 2.73 3.64
@@
w/o defense 18.75 55.00 22.50 39.55 38.18 46.37 50.06
ASTRA (Ours) 8.75 25.00 6.25 13.64 9.09 10.00 9.09
Table 5: Transferability in OOD scenarios. We evaluate the transferability of steering vectors derived from the Jailbreak adversarial images with $\epsilon=\frac{16}{255}$ and choose the same $\alpha$ tuned on the Jailbreak validation set. These steering vectors are tested on the MM-SafetyBench [25] (structured-based attack) and adversarial images very different from the images used for steering vector construction (perturbation-based attack).
@@ 4.1 Experimental Setup
Evaluation Datasets. We evaluate defense performance under two scenarios: in-distribution (ID) and out-of-distribution (OOD). For the ID scenarios, we choose Toxicity and Jailbreak setups using the perturbation-based attack. We sample 55 benign images from ImageNet [11] and apply the PGD attack [30] to generate 25 and 30 adversarial images for validation, and test sets respectively. The perturbation radius $\epsilon$ is set to $\{\frac{16}{255},\frac{32}{255},\frac{64}{255},\text{unconstrained}\}$. For the OOD scenarios, we evaluate the transferable performance on unseen attacks (i.e., structured-based attacks) using MM-SafetyBench [25]. Additionally, we collect 12 images with distributions differing from images used for steering vector construction (e.g., stripes, sketch, painting, etc) and use the PGD attack to construct the test cases.
As for the textual prompts for the perturbation-based attacks, we choose 50 and 100 queries from RealToxicityPrompt [13] to construct the validation and test set for Toxicity setup. We choose 110 and 110 queries from Advbench [58] and Anthropic-HHH [12] to construct the validation and test set for Jailbreak setup. All test queries are apart from the question-answer pairs used in the PGD attack. During the evaluation, we pair each textual prompt with a random adversarial image.
For the evaluation of utility performance in benign scenarios, we employ two established benchmarks, MM-Vet [52] and MM-Bench [28]. Full Details of dataset statistics can be found in Appendix 7.1.
Evaluation Metrics. For the toxicity setup, we follow Qi et al. [37] and use the Detoxify classifier [18] to calculate the toxicity score. We report the average scores of Toxicity attribute across the test set. The scores range from 0 (least toxic) to 1 (most toxic). For the jailbreak setup, we choose the classifier from HarmBench [31] to compute the attack success rate (ASR) for both perturbation-based and structured-based jailbreak.
@@ 4.2 Defense Performance Comparision (RQ1)
MM-Vet [52] MMBench [28] Toxicity (Perturbation-based) Jailbreak (Perturbation-based) Jailbreak (Structured-based)
Direct ASTRA Direct ASTRA Direct ASTRA Direct ASTRA Direct ASTRA
MiniGPT-4 19.40 20.62 35.90 35.82 51.42 10.14 3.95 5.82 2.62 4.29
LLaVA-v1.5 32.62 30.55 72.94 73.23 63.68 59.28 3.68 8.59 3.82 4.61
Qwen2-VL 49.13 48.66 78.00 78.69 140.44 40.14 6.80 8.86 30.00 30.92
@@
Transferability. In real-world settings, unknown distributions and types of adversarial images highlight the need for a robust and transferable defense framework. To evaluate transferability of our steering vectors, we construct two testing scenarios: in-distribution (ID) and OOD cases.
In the ID scenario, we assess whether steering vectors derived from adversarial images with a specific $\epsilon$ value can defend against adversarial images with varying $\epsilon$ levels. Adversarial images used for steering vector construction and test evaluations are drawn from distinct classes in ImageNet [11], ensuring similar image distributions. As illustrated in Fig. 3 and 6, the results demonstrate the effectiveness of our steering vectors defending against adversarial attacks with different $\epsilon$ values. We also report the Avg. performance, in which we take the average of steering vectors derived from adversarial images with $\epsilon$ values in {16, 32, 64, unconstrained}. Despite that the defense with $\epsilon=\text{unconstrained}$ does not work quite well against perturbation-based attacks with $\epsilon=\{\frac{16}{255},\frac{32}{255},\frac{64}{255}\}$, remaining defense validate the transferability of ASTRA across PGD attacks with different intensities.
In the OOD scenarios, we test whether steering vectors derived from the Jailbreak adversarial images with $\epsilon=\frac{16}{255}$ can generalize to different attack types. Specifically, for the structure-based attack, we construct an OOD benchmarking using MM-SafetyBench [25]; for the perturbation-based attack, we collect 12 images from the Internet with characteristics (e.g., stripes, sketch, painting, etc) highly different from the adversarial images used for steering vector construction. Results in Table 5 confirm our framework’s efficacy in both structured and perturbation-based OOD setups, indicating great potential for real-world deployment. This impressive OOD transferability may arise from the steering vectors encapsulating a common harmful feature direction that persists regardless of how the harmful behavior is triggered. Although models can be jailbroken by different types of attacks, eventually, there exists a certain direction in the feature space that represents the harmfulness. By accurately steering away from this direction, we can effectively safeguard models against diverse types of jailbreaks.
@@ 4.3 General Utility (RQ2)
Utility Performance. We calculate utility scores in MM-Vet [52] and MMBench [28] datasets for benign scenario evaluation and perplexity for adversarial scenario evaluation. See Appendix 7.1 for detailed descriptions of utility scores. As shown in Table 6, our defended models demonstrate considerable utility performance in benign scenarios compared to those without defenses. These comparisons demonstrate that our defense results in little performance drops on benign inputs. We owe these results to our adaptive steering approach, which mitigates utility degradation by computing the projection between the language model’s calibrated activation and steering vectors, thereby avoiding the drawbacks of a fixed steering coefficient. In adversarial contexts, the perplexities of ASTRA are still within a reasonable range, indicating that our defended models consistently provide valid, non-harmful responses to harmful instructions. Additional cases are provided in Appendix 7.4.
@@ 7.1 Dataset Statistics
We use the PGD attack to inject adversarial noise into each benign image. For the Jailbreak setup, we prepare 416 and 415 harmful instructions and corresponding submission responses from the AdvBench [58] and Anthropic-HHH [12] respectively to conduct the PGD attack. For the Toxicity setup, we choose 66 toxic queries from Qi et al. [37] as the optimization objective to conduct the PGD attack. We apply PGD for 2500 iterations with a step size of 1/255 on MiniGPT-4, and a step size of 1/1020 on Qwen2-VL and LLaVA-v1.5.
@@ 7.4 Qualitative Results
Qualitative results for Qwen2-VL, LLaVA-v1.5, and MiniGPT-4 under adversarial scenarios are shown in Fig. 910, and 11 respectively, while results under benign scenarios are provided in Fig. 12.
@@ 7.6 Implementation Details
Table 9: Hyperparameter selection of steering coefficient $\alpha$.

### v3

@@ Abstract
Vision Language Models (VLMs) can produce unintended and harmful content when exposed to adversarial attacks, particularly because their vision capabilities create new vulnerabilities. Existing defenses, such as input preprocessing, adversarial training, and response evaluation-based methods, are often impractical for real-world deployment due to their high costs. To address this challenge, we propose ASTRA, an efficient and effective defense by adaptively steering models away from adversarial feature directions to resist VLM attacks. Our key procedures involve finding transferable steering vectors representing the direction of harmful response and applying adaptive activation steering to remove these directions at inference time. To create effective steering vectors, we randomly ablate the visual tokens from the adversarial images and identify those most strongly associated with jailbreaks. These tokens are then used to construct steering vectors. During inference, we perform the adaptive steering method that involves the projection between the steering vectors and calibrated activation, resulting in little performance drops on benign inputs while strongly avoiding harmful outputs under adversarial inputs. Extensive experiments across multiple models and baselines demonstrate our state-of-the-art performance and high efficiency in mitigating jailbreak risks. Additionally, ASTRA exhibits good transferability, defending against unseen attacks (i.e., structured-based attack, perturbation-based attack with project gradient descent variants, and text-only attack). Our code is available at https://github.com/ASTRAL-Group/ASTRA.
@@ 1 Introduction
Jailbreak attacks in VLMs aim to induce models to generate harmful responses by using jailbreaking image-text pairs [24, 23, 22, 51, 48, 41, 26]. These jailbreak attacks can be categorized into two types: (i) perturbation-based attacks, which create adversarial images that prompt generation of the harmful response from VLMs [41, 46, 2, 38], (ii) structured-based attacks, which embeds the malicious queries into images via typography to bypass the safety alignment of VLMs [17, 30]. Countermeasures for both attacks have been explored extensively: the input preprocessing-based method [37] or adversarial training [25] have proven effective for perturbation-based attacks. However, these defenses suffer as they require intensive computations to purify the image or fine-tune the model. Response evaluation-based  [18, 53, 58] defenses have been proposed for structured-based attacks, but they all require running model inference multiple times to potentially identify harmful outputs, which dramatically increases the cost of real-world deployment.
@@
    ASTRA achieves a substantial improvement in defending against perturbation-based attacks. Compared to state-of-the-art methods JailGuard [58], with a Toxicity Score of 12.12% and an Attack Success Rate of 17.84% lower, and 9x faster in MiniGPT-4. ASTRA is also transferable to some unseen attacks (i.e., structure-based attack, perturbation-based attack with PGD variants, and text-only attack), and still be effective against adaptive attacks.
@@ 2 Related Work
Jailbreak Attacks on VLM. Jailbreak attacks aim to alter the prompt to trick the model into answering forbidden questions. Apart from the LLM-based textual jailbreak strategies [31, 64, 19, 56, 60], additional visual inputs expose a new attack surface to VLM attacks. There are two main types of attacks: perturbation-based attacks and structured-based attacks [53]. Perturbation-based attacks create adversarial images to bypass the safeguard of VLMs [6, 41, 2, 46, 55, 59]. Structued-based attacks convert the harmful content into images through typography or text-to-image tool (e.g., Stable Diffusion [44]) to induce harmful responses from the model [29, 17, 30, 33, 27] We study our defense on both types of attacks.
@@ 3 Methodology
         Compute: $f(g_{n})=\text{log}\mathcal{P}_{\text{VLM}}(\mathbf{r}|\text{Ablate}(\mathbf{x}_{v},g_{n}),\mathbf{x}_{t})$
@@ 3.1 Constructing Steering Vectors
Harmful Feature Extraction. With attribution scores for each token, we extract the representation of those tokens strongly correlated with jailbreak. Additionally, we hope our steering vectors generalize rather than overfitting to specific instructions and enjoy good transferability to a wider range of jailbreaks. Thus, we utilize visual tokens with Top-$k$ attribution scores from the surrogate model $\hat{f}$ paired with the empty user query to construct the steering vectors.
@@
$v^{l}=\frac{1}{|\mathcal{D}|}\sum_{\mathbf{x}_{v},\text{Mask}(\mathbf{x}_{v})\in\mathcal{D}}$ $\left[\mathbf{a}^{l}\left(\mathbf{x}_{v},\mathbf{x}_{\text{template}}\right)\right.$
$\left.-\mathbf{a}^{l}\left(\text{Mask}(\mathbf{x}_{v}),\mathbf{x}_{\text{template}}\right)\right]$
@@ 4 Experiments
Table 5: Transferability in OOD scenarios. We evaluate the transferability of steering vectors derived from the Jailbreak adversarial images with $\epsilon=\frac{16}{255}$ and choose the same $\alpha$ tuned on the Jailbreak validation set. The transferability is evaluated across multiple unseen attack categories: structured-based attack from MM-SafetyBench [29], perturbation-based attack with various PGD variants, and text-only attack. We use the classifier from HarmBench [35] to compute the ASR.
Structured-based Attack Perturbation-based Attack Text-only Attack
SD SD_TYPO TYPO PGD [34] Auto-PGD [11] MI-FGSM [14] GCG [64]
$\epsilon=16/255$ $\epsilon=32/255$ $\epsilon=16/255$ $\epsilon=32/255$ $\epsilon=16/255$ $\epsilon=32/255$
@@
w/o defense 13.75 43.25 43.75 70.91 78.18 74.55 76.36 78.18 79.09 58.18
ASTRA (Ours) 3.75 8.75 11.25 5.45 12.73 5.45 10.91 16.37 13.64 9.09
@@
w/o defense 20.00 61.25 38.75 74.55 80.00 76.37 77.57 80.00 78.18 81.82
ASTRA (Ours) 11.25 40.00 33.75 21.82 14.55 15.76 15.76 18.18 18.18 30.91
@@
w/o defense 18.75 55.00 22.50 69.09 74.55 80.60 90.30 87.28 89.09 92.73
ASTRA (Ours) 8.75 25.00 6.25 1.82 1.82 1.21 0.61 0.00 0.00 14.55
@@ 4.1 Experimental Setup
Evaluation Datasets. We choose Toxicity and Jailbreak setups using the perturbation-based attack. We sample 55 benign images from ImageNet [13] and apply the PGD attack [34] to generate 25 and 30 adversarial images for visual validation, and test sets respectively. The perturbation radius $\epsilon$ is set to $\{\frac{16}{255},\frac{32}{255},\frac{64}{255},\text{unconstrained}\}$. For textual prompts, we choose 50 and 100 queries from RealToxicityPrompt [16] to construct the validation and test set for Toxicity setup. We choose 110 and 110 queries from both Advbench [64] and Anthropic-HHH [15] to construct the validation and test set for Jailbreak setup. All text prompts are different from the instruction-response pairs used for steering vector construction. During the evaluation, we pair each textual prompt with a random adversarial image. Detailed data information can be found in Fig. 4.
For the evaluation of utility performance in benign scenarios, we employ two established benchmarks, MM-Vet [57] and MM-Bench [32]. Additionally, we include safe instructions from XSTest [45] to assess the overrefusal case. Full Details of dataset statistics can be found in Appendix 8.1.
Evaluation Metrics. For the toxicity setup, we follow Qi et al. [41] and use the Detoxify classifier [21] to calculate the toxicity score. We report the average scores of Toxicity attribute across the test set. The scores range from 0 (least toxic) to 1 (most toxic). For the jailbreak setup, we choose the classifier from HarmBench [35] to compute the attack success rate (ASR).
@@ 4.2 Defense Performance Comparision (RQ1)
MM-Vet [57] MMBench [32] XSTest [45] Toxicity (Perturbation-based) Jailbreak (Perturbation-based) Jailbreak (Structured-based)
Direct ASTRA Direct ASTRA Direct ASTRA Direct ASTRA Direct ASTRA Direct ASTRA
MiniGPT-4 19.40 20.62 35.90 35.82 87.60 87.60 51.42 10.14 3.95 5.82 2.62 4.29
LLaVA-v1.5 32.62 30.55 72.94 73.23 98.00 98.80 63.68 59.28 3.68 8.59 3.82 4.61
Qwen2-VL 49.13 48.66 78.00 78.69 73.60 74.00 140.44 40.14 6.80 8.86 30.00 30.92
@@
Transferability. In real-world scenario, unknown types of adversarial attacks highlight the need for a robust and transferable defense framework. To evaluate transferability of ASTRA, we construct two test cases: in-distribution (ID) and out-of-distribution (OOD).
In ID scenario, adversarial images used for steering vector construction and test evaluations are drawn from same classes in ImageNet [13], ensuring similar image distributions. We assess whether steering vectors derived from adversarial images with a specific $\epsilon$ value can defend against adversarial images with varying $\epsilon$ levels. As illustrated in Fig. 3 and 7, the results demonstrate the effectiveness of our steering vectors defending against adversarial attacks with different $\epsilon$ values. We also report the Avg. performance, in which we take the mean of steering vectors derived from adversarial images with $\epsilon$ values in $\{\frac{16}{255},\frac{32}{255},\frac{64}{255},\text{unconstrained}\}$. Despite that the defense with $\epsilon=\text{unconstrained}$ does not work quite well against perturbation-based attacks with $\epsilon=\{\frac{16}{255},\frac{32}{255},\frac{64}{255}\}$, remaining defense validate the transferability of ASTRA across PGD attacks with different intensities.
In OOD scenario, we test whether steering vectors derived from the Jailbreak adversarial images with $\epsilon=\frac{16}{255}$ can generalize to different types of attacks. Specially, we evaluate the defense transability on structured-based attack from MM-SafetyBench [29], perturbation-based attack with several PGD variants, and text-only attack. Please refer Appendix 8.1 for details of structured-based attack. For the perturbation-based attack, we collect 12 images with distributions differing from images used for steering vector construction (e.g., stripes, sketch, painting, etc). We use 55 instruction-response pairs from JailbreakBench [7] to conduct perturbation-based attacks with PGD variants (i.e, PGD, MI-FGSM [14], and Auto-PGD [11]) and text-only attack (i.e., GCG [64]). We use the same 55 instructions from JailbreakBench [7] to evaluate performance.
Results in Table 5 confirm the defense transferability across different unseen attacks, indicating great potential for real-world deployment. This impressive OOD transferability may arise from the steering vectors encapsulating a common harmful feature direction that persists regardless of how the harmful behavior is triggered. Although models can be jailbroken by different types of attacks, eventually, there exists a certain direction in the feature space that represents the harmfulness. By accurately steering away from this direction, we can effectively safeguard models against diverse types of jailbreaks.
@@ 4.3 General Utility (RQ2)
Utility Performance. We calculate utility scores in MM-Vet [57], MMBench [32], and safe instructions from XSTest [45] for benign scenario evaluation and perplexity for adversarial scenario evaluation. See Appendix 8.1 for detailed descriptions of utility scores. As shown in Table 6, our defended models demonstrate considerable utility performance in benign scenarios compared to those without defenses. These comparisons demonstrate that our defense results in little performance drops on benign inputs. We owe these results to our adaptive steering approach, which mitigates utility degradation by computing the projection between the language model’s calibrated activation and steering vectors, thereby avoiding the drawbacks of a fixed steering coefficient. In adversarial contexts, the perplexities of ASTRA are still within a reasonable range, indicating that our defended models consistently provide valid, non-harmful responses to harmful instructions. Additional cases are provided in Appendix 8.4.
@@ 5 Conclusion
6 Acknowledgment
This work was supported by NSF 2331967, 2229876 and 2055233. Huan Zhang is supported in part by the AI2050 program (Early Career Fellowship) at Schmidt Sciences.
@@ 6 Acknowledgment
Figure 4: Data used for steering vectors construction and test evaluation in the Jailbreak setup.
@@ 8.1 Dataset Statistics
We use the PGD attack to inject adversarial noise into each benign image. For the Jailbreak setup, we prepare 416 and 415 harmful instructions and corresponding affirmation responses from the AdvBench [64] and Anthropic-HHH [15] respectively to conduct the PGD attack. For the Toxicity setup, we choose 66 toxic queries from Qi et al. [41] as the optimization objective to conduct the PGD attack. We apply PGD for 2500 iterations with a step size of 1/255 on MiniGPT-4, and a step size of 1/1020 on Qwen2-VL and LLaVA-v1.5.
@@
XSTest [45] evaluates the exaggerated safety behavior in LLMs. To construct our test set, we use 250 safe prompts across ten prompt types that well-calibrated models should not refuse to comply with. We define the utility score as the proportion of safe prompts with which the model complies, measured via the string matching.
@@ 8.4 Qualitative Results
Qualitative results for Qwen2-VL, LLaVA-v1.5, and MiniGPT-4 under adversarial scenarios are shown in Fig. 1011, and 12 respectively, while results under benign scenarios are provided in Fig. 13.
@@ 8.6 Implementation Details
Table 9: Hyperparameter selection of steering coefficient $\alpha$.